If 27 and 45 are two lengths of a right triangle, what is the third length of the right triangle?
Answer by Edwin McCravy(20055)

If 27 and 45 are two lengths of a right triangle, what is the third length of the right triangle?
There are two possible answers because you didn't say which one was 
the hypotenuse:

Case 1:  27 and 45 are both legs and the third side is the hypotenuse:

        c² = a² + b²
        c² = 27² + 45²
        c² = 729 + 2025
        c² = 2754
         c = 52.47856705 (taking the square root)

 
Case 2: 45 is the hypotenuse, 27 is one of the legs, and the third 
side is the other leg:
        c² = a² + b²
       45² = 27² + b²
      2025 = 729 + b²
2025 - 729 = b²
      1296 = b²
        36 = b  (taking the square root)

Edwin